(4/5)x + 18 = 12
(4/5)x = 12 + 18
(4/5)x = 30
x = 30(4/5)

Answer by fractalier(2101)

That isn't quite right...
(4/5)x + 18 = 12
(4/5)x = 12 - 18
(4/5)x = -6
x = (-6)(5/4)
x = -30/4 = -15/2
is correct...